Explain PVA vs FVA and give examples.
Answer #1
PVA:
PVA means PV of annuity. Annuity is series
of equal payments for a specified number
of periods.
The PV of annuity = [(1+r)^n-1]/[r*(1+r)^n]
I needed an annuity of $500 (yearly) for 25 years
for which I approached an Insurance co.
The company asked me to pay the PV of the
annuity, the discount rate being 3% p.a.
The PV of annuity or the amount payable today to purchase the annuity = 500*(1.03^25-1)/(0.03*1.03^25) = $         8,707
FVA is the opposite of PVA.
It is given by the formula:
FVA = [(1+r)^n-1]/r
Example:
I plan to save $500 for 5 years at the end of each
year. If I can earn an interst rate of 3% compounded
annually, the FV of this annuity will be:
500*(1.03^5-1)/0.03 = $         2,655